Suppose that a pair of dice is tossed. One die has 9 sides, the other has 7 sides. What is the expected value of the sum of the two dice? (Don't write out the full probability model, add the two expected values.)
Answer:

Answer by greenestamps(13198)



Obviously the question as posted can't be answered, because there is no information about what numbers are on the faces of each die....

Assuming the numbers are the whole numbers from 1 to n for a die with n faces, the expected value for any roll on each die is (1+n)/2.

Then the expected value on a roll of the 7-sided die is (1+7)/2 = 4 and the expected value on a roll of the 9-sided die is (1+9)/2 = 5; so the expected value of the sum of one roll on each die is 4+5 = 9.

ANSWER: 9